Find dy/dx and d²y/dx² without eliminating the parameter.


a.) x= e^(2t) , y= t In(t)

b.) x= sinh(t) , y= cosh(t)

c.) x= 1-t^2 , y= 1+t


1
Expert's answer
2021-01-13T19:28:24-0500

We shall use the formulas:


"\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}},\ \ \frac{d^2y}{dx^2}=\frac{\frac{d}{dt}(\frac{dy}{dx})}{\frac{dx}{dt}}"


a.) "x= e^{2t} , y= t\ln(t)"


"\frac{dy}{dx}=\frac{\ln(t)+t\frac{1}{t}}{2e^{2t}}=\frac{\ln(t)+1}{2e^{2t}}"


"\frac{d^2y}{dx^2}=\frac{1}{2e^{2t}}\cdot \frac{\frac{1}{t}2e^{2t}-(\ln(t)+1)4e^{2t}}{4e^{4t}}=
 \frac{1-2t(\ln(t)+1)}{4te^{4t}}"



b.) "x= \sinh(t) , y= \cosh(t)"


"\frac{dy}{dx}=\frac{\sinh(t)}{\cosh(t)}=\tanh(t)"


"\frac{d^2y}{dx^2}=\frac{1-\tanh^2(t)}{\cosh(t)}"



c.) "x= 1-t^2 , y= 1+t"


"\frac{dy}{dx}=\frac{1}{-2t}=-\frac{1}{2t}"


"\frac{d^2y}{dx^2}=\frac{1}{-2t}\cdot\frac{1}{2t^2}=-\frac{1}{4t^3}"